OpenStudy (anonymous):

Find the zeros of the function.f(x) = 9 x3 - 72 x2 + 135x

OpenStudy (callisto):

Put f(x) = 0 \[9x^3 - 72x^2 + 135x =0\]\[9x(x^2 - 8x + 15) =0\]\[9x(x-3)(x-5) =0\]Put each factor = 0 and solve for x.